solve using the square root property: 3(x-4)^2=48

Answer by sudhanshu_kmr(1152)


3(x-4)^2=48
=> (x-4)^2 = 48/3
=> (x-4)^2 = 16
so, x-4 = 4 or x-4 = -4

x = 8 or x = 0